### v1

# Nonparametric Jackknife Instrumental Variable Estimation and Confounding Robust Surrogate Indices

Abstract
Jackknife instrumental variable estimation (JIVE) is a classic method to leverage many weak instrumental variables (IVs) to estimate linear structural models, overcoming the bias of standard methods like two-stage least squares. In this paper, we extend the jackknife approach to nonparametric IV (NPIV) models with many weak IVs. Since NPIV characterizes the structural regression as having residuals projected onto the IV being zero, existing approaches minimize an estimate of the average squared projected residuals, but their estimates are biased under many weak IVs. We introduce an IV splitting device inspired by JIVE to remove this bias, and by carefully studying this split-IV empirical process we establish learning rates that depend on generic complexity measures of the nonparametric hypothesis class. We then turn to leveraging this for semiparametric inference on average treatment effects (ATEs) on unobserved long-term outcomes predicted from short-term surrogates, using historical experiments as IVs to learn this nonparametric predictive relationship even in the presence of confounding between short- and long-term observations. Using split-IV estimates of a debiasing nuisance, we develop asymptotically normal estimates for predicted ATEs, enabling inference.
1 Introduction
The non-parametric instrumental variable (NPIV) problem is, given $N$ observations, $O_{i}=(A_{i},S_{i},Y_{i})$ $i=1,\dots,N$, on instrumental variables (IVs) $A$, interventions $S\in\mathbb{R}^{p}$, and outcomes $Y\in\mathbb{R}$, to find a function $h:\mathbb{R}^{p}\to\mathbb{R}$ satisfying
$\mathbb{E}[Y-h(S)\mid A]=0.$ (1)
This can be motivated by a structural (i.e., causal) model $Y=h^{\star}(S)+\epsilon$, where $S$ and $\epsilon$ can be endogenous due, e.g., to the presence of common confounders (so that $h^{\star}(S)\neq\mathbb{E}[Y\mid S]$), but $A$ and $\epsilon$ are exogenous so that Eq. 1 holds for $h^{\star}$. A variety of work studies the estimation of $h^{\star}$ and inference on functionals thereof under nonparametric restrictions on $h^{\star}$ as we receive additional observations $N$ from a fixed $O=(A,S,Y)$ distribution (Newey and Powell, 2003; Severini and Tripathi, 2006; Ai and Chen, 2003, 2007, 2012; Chen and Pouzo, 2009; Severini and Tripathi, 2012; Bennett and Kallus, 2023; Bennett et al., 2019, 2023a; Chen and Pouzo, 2012; Darolles et al., 2011; Dikkala et al., 2020; Hartford et al., 2017; Santos, 2011; Singh et al., 2019; Zhang et al., 2023; Kremer et al., 2022, 2023; Bennett et al., 2023b, b).
In this paper, we tackle NPIV in the challenging many-weak-IV setting, where $A\in\{1,\dots,K\}$ is discrete and we only see so many of each value, namely for each $a\in[K]$, we have $n$ i.i.d. observations of $(Y,S)\mid A=a$, forming $N=nK$ observations in total.
In the linear setting where $h^{\star}(S)=S^{\top}\beta^{\star}$, Eq. 1 reduces to solving $\mathbb{E}[Y\mid A]=\mathbb{E}[S\mid A]^{\top}\beta$ for $\beta$. This motivates the two-stage least squares (2SLS) approach of estimating $\beta$ by ordinary least squares (OLS) of $Y$ on the “first-stage" OLS prediction of $S$ given $A$ (for discrete $A$ this is simply the sample means of $S$ for each $A$ value). However, when $n\ll N$, even as $N\to\infty$ this can incur non-vanishing bias because the first-stage regression may not converge at all (Angrist et al., 1999; Peysakhovich and Eckles, 2018; Bibaut et al., 2024). JIVE (Angrist et al., 1999) addresses this by regressing $Y$ on a prediction of $S$ given $A$ based on OLS using all the data except the datapoint on which we make the prediction. This renders the errors from the first stage uncorrelated with the second stage so they average out to zero so that we regain consistency (Chao et al., 2012).
The NPIV analog, which we tackle, is, however, unresolved. It is also rather nuanced because when $S$ is continuous but $A$ is discrete, a nonparametric $h^{\star}$ is generally not uniquely identified by Eq. 1, which involves just $K$ moments but a general function $h$. Nonetheless, certain linear functionals of $h^{\star}$, meaning $\theta_{0}=\mathbb{E}[\alpha(S)h^{\star}(S)]$ for some $\alpha$, may still be uniquely identified, meaning $\theta_{0}=\mathbb{E}[\alpha(S)h(S)]$ for any $h$ satisfying Eq. 1.
This problem setting is of particular interest in digital experimentation, where the rapid pace of innovation means we have many ($K$) historical randomized experiments (with serial numbers $A$), which can be used to instrument for the effect ($h^{\star}$) of short-term surrogate observations ($S$) on long-term outcomes ($Y$) even in the presence of unobserved confounding between the two, but where each experiment has a certain sample size ($n$). If we know this effect, we can construct a surrogate index $h^{\star}(S)$ such that average treatment effects (ATEs) on $Y$ are the same as those on $h^{\star}(S)$. Moreover, the ATEs on $h^{\star}(S)$ is a linear functional thereof. Then, for novel experiments, predicting long-term ATEs before observing long-term outcomes can be phrased as inference on a linear functional of a solution to Eq. 1.
In this paper we also tackle the question how to reliably do this inference in the presence of underidentified and nonparameteric $h^{\star}$, which is another significant challenge, besides solving Eq. 1 in the many-weak-IV setting. We furthermore extend the simple instrumentation identification to account for the possibility that short-term surrogate observations do not fully mediate the treatment effects on long-term outcomes (that is, there is exclusion violation).
In this paper, we develop both a novel estimator for $h^{\star}$ in the nonparametric many-weak-IV setting and methods for debiased inference on surrogate-predicted ATEs. The contributions and organization of the paper are as follows
  1. 1.
    In Section 2, we propose a novel nonparametric jackknife IV estimator (npJIVE) based on estimating the average of squared moments using a split-IV device inspired by JIVE and then minimizing it over a generic nonparametric hypothesis class with Tikhonov regularization.
  2. 2.
    In Section 2.1, we prove that npJIVE is consistent to the minimum norm function satisfying Eq. 1 and that its average of squared moments converges at a rate governed by the functional complexity of the hypothesis. To our knowledge, these guarantees are the first of their kinds for NPIV in the many-weak-IV setting.
  3. 3.
    In Section 3, we study the above IV-surrogate-index approach to long-term causal inference and extend it to also tackle the setting where $S$ does not fully mediate all of the effect on $Y$. We establish identification conditions where we can combine historical experiments with full long-term observations together with a new experiment with only short-term observations in order to identify the long-term causal effect in the new experiment, even in the presence of confounding and exclusion violations.
  4. 4.
    In Section 4, we a develop methods for inference on the new experiment’s long-term causal effect. We devise a debiased estimator that involves a new debiasing nuisance related to the ratio of densities of short-term outcomes in the old and the new experiments. We develop an estimator for this debiasing nuisance with guarantees in the many-weak-IV setting. And, we show that when we combine npJIVE with this debiasing nuisance in the right way, we can obtain asymptotically normal estimates of the long-term causal effect in the new experiment, even when npJIVE and the debiasing nuisance are learned completely nonparametrically. This ensures fast estimation rates and implies simple Wald confidence intervals give asymptotically correct coverage.
Taken together, our methods and results provide new ways to conduct long-term causal inference in challenging, but practically very relevant, settings.
2 The npJIVE Estimator for NPIV with Many Weak IVs
We now motivate and define npJIVE. For function $f(A,S,Y)$ let us write $[\Pi f](A)=\mathbb{E}[f(A,S,Y)\mid A]$ and $\|f\|^{2}=\frac{1}{K}\sum_{a=1}^{K}\|f\|^{2}_{(a)}$ with $\|f\|^{2}_{(a)}=\mathbb{E}[f^{2}(a,S,Y)\mid A=a]$, and let $L_{2}(S)=\{h(S):\|h\|<\infty\}$ be the square integrable functions of $S$. Let
$\mathcal{H}_{0}=\{h\in L_{2}(S):R_{1}(h)=0\},~{}~{}\text{where}~{}~{}R_{1}(h)=\frac{1}{2}\|\Pi(h-Y)\|^{2},$
be the solutions to Eq. 1. Given a hypothesis class $\mathcal{H}\subseteq L_{2}(S)$, suppose $\mathcal{H}$ contains a solution $h_{0}\in\mathcal{H}\cap\mathcal{H}_{0}\neq\varnothing$. Then $h_{0}\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}R_{1}(h)\subseteq\mathcal{H}_{0}$, meaning the solution must minimize the risk criterion $R_{1}(h)$ over $\mathcal{H}$ and any such minimizer must solve Eq. 1.
Then, an approach to recover such a solution from data would be to estimate $R_{1}$ and minimize it. Perhaps the most natural empirical analog of $R_{1}$ is the empirical “plug-in” risk
$\widehat{R}_{1}^{\mathrm{plug-in}}(h)=\frac{1}{2K}\sum_{a=1}^{K}([\widehat{\Pi}(Y-h)](a))^{2},~{}~{}\text{where}~{}~{}[\widehat{\Pi}f](a)=\frac{1}{n}\sum_{i:A_{i}=a}f(a,S_{i},Y_{i}),$ (2)
meaning $[\widehat{\Pi}f](a)$ is the sample average over the $A_{i}=a$ data. In fact, optimizing $\widehat{R}_{1}^{\mathrm{plug-in}}(h)$ over linear functions $h(S)=S^{\top}\beta$ is almost 2SLS (aside from some $A$-dependent sample weights). And, optimizing it over a nonparametric hypothesis class $\mathcal{H}$ (possibly with regularization) is exactly the adversarial NPIV estimators of (Bennett et al., 2019, 2023a, 2023b; Bennett and Kallus, 2023; Dikkala et al., 2020) with an adversary function class that includes all functions of the discrete $A$.
Unfortunately, this is not an unbiased estimate of $R_{1}(h)$, and it can be inconsistent when $n\not\to\infty$. Essentially, given $n$ observations on some variables $(Z,Z^{\prime})$, the expectation of the product of sample means of $Z_{i}$ and of $Z_{i}^{\prime}$ is not the product of the expectations of $Z$ and $Z^{\prime}$ – there’s also $1/n$ times the covariance of $Z$ and $Z^{\prime}$. This is the fundamental reason why 2SLS isn’t consistent under many weak IVs. One way to see the key trick in JIVE is that, partitioning observations in folds $1,\ldots,L$ of equal size, with $V_{i}\in[L]$ the fold membership of observation $i$, an unbiased estimate of the product of expectations is the product of the empirical mean of $Z_{i}$ over folds $11,\ldots,L-$ 1 with the empirical mean of $Z_{i}^{\prime}$ over fold $L$. (JIVE uses $L=n$, corresponding to “leave one out" sample means).
Drawing inspiration from this insight, we propose a 2-fold cross-fold risk estimator:
$\widehat{R}_{1}(h)$ $=\frac{1}{2K}\sum_{a=1}^{K}[\widehat{\Pi}_{0}(Y-h)](a)[\widehat{\Pi}_{1}(Y-h)](a),$
$\text{where}~{}~{}[\widehat{\Pi}_{v}f](a)$ $=\frac{2}{n}\sum_{i:A_{i}=a,V_{i}=v}f(a,S_{i},Y_{i}),$
assuming $n$ is even. Meaning $[\widehat{\Pi}_{v}f](a)$ takes the sample average over the observations for which $A_{i}=a$ belonging to fold $v$.
Our proposed npJIVE estimator is to minimize this risk plus a Tikhonov regularization penalty
$\hat{h}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}}~{}\widehat{R}_{1}(h)+\lambda\|h\|_{2,N}^{2},$ (npJIVE)
where $\|h\|_{2,N}^{2}=\frac{1}{N}\sum_{i=1}^{N}h^{2}(S_{i})$. In fact, optimizing $\widehat{R}_{1}(h)$ over linear functions $h(S)=S^{\top}\beta$ is almost JIVE (aside from leave-one-out instead of 2 folds and some $A$-dependent sample weights).
2.1 Analysis of npJIVE
We now give an analysis of the convergence of npJIVE, both in terms of violations of Eq. 1 and in terms of distance to a minimum norm solution.
First we need to assume some boundedness. Henceforth, let us assume throughout that $\left\lvert Y\right\rvert\leq 1$ and that $\|h\|_{\infty}\leq 1$ for every $h\in\mathcal{H}$.
Second we need to characterize the complexity of our function class $\mathcal{H}$. Let $\mathcal{D}_{v}$ be the set of observations belonging to fold $v$, that is $\mathcal{D}_{v}=\{O_{i}:i\in[N],~{}V_{i}=v\}$. Let $\epsilon_{1},\ldots,\epsilon_{N}$ be i.i.d. Rademacher random variables independent from $O_{1},\ldots,O_{N}$. For any function class $\mathcal{F}$, we introduce the following Rademacher complexities:
$\mathcal{R}_{N}(\mathcal{F},\delta)=$ $\mathbb{E}\left[\sup_{f\in\mathcal{F}:\|f\|\leq\delta}\frac{2}{N}\sum_{i:V_{i}=0}\epsilon_{i}f(S_{i})\right],$ (3)
$\mathcal{R}_{n,a}(\mathcal{F},\delta)=$ $\mathbb{E}\left[\sup_{f\in\mathcal{F}:\|f\|_{(a)}\leq\delta}\frac{2}{n}\sum_{i:A_{i}=a,V_{i}=0}\epsilon_{i}f(S_{i})\right],$ (4)
$\text{and}\qquad\mathcal{R}_{N}^{|0}(\mathcal{F},\delta)=$ $\mathbb{E}\left[\sup_{f\in\mathcal{F}:\|f\|\leq\delta}\frac{2}{N}\sum_{i:V_{i}=1}\epsilon_{i}f(S_{i})\mid\mathcal{D}_{0}\right].$ (5)
The relevant function classes for analyzing npJIVE are: $\mathcal{H}^{0}=\mathcal{H}-h_{0}$, $\mathcal{G}^{0}=\{g(S)-(\Pi g)(A):g\in\mathcal{H}^{0}\}$, $\mathcal{G}^{1}=\{g(S)(\Pi g)(A):g\in\mathcal{H}^{0}\}$, and $\widehat{\mathcal{G}}=\{g(S)(\widehat{\Pi}_{0}g)(A):g\in\mathcal{G}^{0}\}$.
Let $\mathrm{star}(\mathcal{F})=\{\gamma g:\gamma\in[0,1],g\in\mathcal{F}\}$. Let $\widetilde{\delta}_{N}$, $\bar{\delta}_{N}$ and $\breve{\delta}_{N}$ be solutions to $\mathcal{R}_{N}(\mathrm{star}(\mathcal{F}),\delta)\leq\delta^{2}$ for $\mathcal{F}=\mathcal{G}^{0},\mathcal{G}^{1},\Pi\mathcal{G}$, respectively, let $\widetilde{\delta}_{n}$ be a solution to $\mathcal{R}_{n,a}(\mathrm{star}(\mathcal{G}^{0}),\delta)\leq\delta^{2}$ for every $a$, and let $\widehat{\delta}_{N}$ be a solution to $\mathcal{R}_{N}^{|0}(\mathrm{star}(\widehat{\mathcal{G}}),\delta)\leq\delta^{2}$. Let $\widehat{r}_{N}^{2}$ and $\bar{r}_{N}^{2}$ be bounds on $\mathcal{R}^{|0}(\widehat{\mathcal{G}},\infty)$ and $\mathcal{R}(\mathcal{G}^{1},\infty)$.
Here, $\widehat{r}_{N}^{2},\bar{r}_{N}^{2}$ are (unlocalized) Rademacher complexity measures and $\widetilde{\delta}_{n},\widetilde{\delta}_{N},\widehat{\delta}_{N},\bar{\delta}_{N},\breve{\delta}_{N}$ are critical radii, both of which are standard measures of generic functional complexity (see (Wainwright, 2019) which also gives bounds on these quantities for a variety of classes including linear, reproducing kernel Hilbert spaces, Hölder or Sobolev spaces, neural nets, etc.). Finally, let $\delta_{N}=\max(\widetilde{\delta}_{N},\widehat{\delta}_{N},\bar{\delta}_{N},\breve{\delta}_{N})$ and $r_{N}^{2}=\max(\widehat{r}_{N}^{2},\bar{r}_{N}^{2})$.
Theorem 1.
Let $\zeta\in(0,0.5)$ be given. Suppose that $\delta_{N}^{2}=o(\lambda)$, $N^{-1/2}\sqrt{\log\log\widehat{\delta}_{N}}=o(\widehat{\delta}_{N})$, $N^{-1/2}\sqrt{\log\log\bar{\delta}_{N}}=o(\bar{\delta}_{N})$, $n^{-1/2}\sqrt{\log\log\widetilde{\delta}_{n}}=o(\widetilde{\delta}_{n})$ and that $\sqrt{\log(K/\zeta)}=o(\widetilde{\delta}_{n})$. Then,
$\|\Pi(\widehat{h}-h_{0})\|^{2}=O(\lambda+\delta_{N}\delta_{n})~{}~{}\text{with probability $1-O(\zeta)$, and}$
$\|\widehat{h}-h_{0}\|=o_{P}(1)$
3 Confounding Robust Surrogate Indices from Many Weak Experiments
We now turn to the problem of using many historical experiments to construct surrogate indices for long-term causal inference in a novel experiment. The long-term effect of interventions is often of primary concern in causal inference. Examples include the effect of early-childhood education on lifetime earnings (Chetty et al., 2011), of promotions on long-term value (Yang et al., 2020), and of digital platform design on long-term user retention (Hohnhold et al., 2015). While the gold standard for causal inference is experimentation, the significant delay of long-term observations after assignment to treatment means that, even when we can randomize the intervention, we may not be able to measure the outcome of interest.
This presents significant challenges and may even alter incentives such that scientists prioritize interventions that can be evaluated with short-run outcomes, as documented by Budish et al. (2015) in the case of cancer drugs. Nevertheless, other relevant post-treatment outcomes are often available in the short-term. For example, in AIDS treatments we observe short-term viral loads or CD4 counts well before mortality outcomes (Fleming et al., 1994). Similarly, in digital experimentation, we observe short-term signals on user engagement well before retention or revenue shifts.
An appealing approach to leverage these short-term observations is in the construction of a surrogate index, whereby we impute unobserved long-term outcome using their prediction from multiple short-term surrogate observations.
3.1 Background: Statistical-Surrogate Indices
Possibly the most common approach assumes that short-term observations form a statistical surrogate (Prentice, 1989). This assumes that the long-term outcome is conditionally independent of treatment given the short-term outcomes. This involves two key restrictions: that there are no unobserved confounding between short- and long-term outcomes (no $U_{1}$ in Fig. 1(a)) and that all of the treatment’s effect on long-term outcomes is mediated by the short term (no $U_{2}$ in Fig. 1(a)). As the latter restriction becomes more defensible as we include more short-term outcomes so as to mediate more of the treatment’s long-term effect, Athey et al. (2019) combine many short-term observations into a surrogate index, assuming they form a statistical surrogate and using historical data to regress long-term on short-term (or other ways of adjustment such as weighting). The proposal is simple, effective, and, as such, widely adopted.
However, even if short-term outcomes fully mediate the long-term effect, they may fail to satisfy statistical surrogacy. Consider the causal diagram in Fig. 1(a), where $S$ perfectly mediates $A$’s effect on $Y$ (i.e., exclusion restriction), but $S$ and $Y$ share an unobserved confounder $U$, while treatment $A$ is fully unconfounded. In this case, $S$ is a collider so that conditioning on it induces a path from $A$ to $Y$ via $U$, violating surrogacy and imperiling analysis using such methods as Athey et al. (2019).
This scenario is actually exceedingly commonplace: for example, in streaming platforms with a subscription model, a user’s amount of free time both impacts their short-term engagement and their probability of subscription retention in the same direction. Failing to address this confounding leads to surrogate-based estimates that overestimate true effects on long-term outcomes. We might even have more extreme situations in which an intervention strongly increases short-term engagement of a subpopulation of users who are unlikely to unsubscribe while slightly decreasing engagement of a subpopulation of users very likely to unsubscribe. This might result in both an overall increase in short-term engagement and an overall decrease in long-term retention, a situation known as the surrogate paradox (Elliott et al., 2015).
3.2 Experiments as Instruments and Surrogates as Proxies
Athey et al. (2019) consider a setting where historical data prior to the present experiment contains only $S$ and $Y$ and some baseline covariates. In this context, all we can do is either worry about potential unobserved confounders, or hope the included covariates satisfy ignorability between $S$ and $Y$. However, at organizations that routinely run many digital experiments, we can take historical data from past experiments where we also observe the randomized treatments $A$. In the setting of Fig. 1(a), these treatments constitute an IV, which can help us identify the causal effect of $S$ on $Y$ and therefore infer the effect of a novel treatment on $Y$ by considering only its effect on $S$.
Going beyond this, we can even account for some exclusion violations (a long-term effect unmediated by short-term observations) under additional structure. In the following we show that some parts of $S$ can serve as effect mediators, while other parts of $S$ can serve as negative controls (also known as proxies) that can adjust for effects unmediated by the first part, so long as we get a rich enough view onto them.
In the following, we first present the corresponding causal model (summarized in Fig. 1(b)). Then, we present an identification result showing how the long-term outcome of a novel treatment can be identified without observations on long-term outcomes under the novel treatment by leveraging historical experiments where we observe both short- and long-term outcomes. This identification result is written as a function of NPIV solutions, which may be estimated slowly in this nonparametric setting, imperiling rates on causal effect estimation if we were to simply plug-in an estimated solution. Therefore, we finally present a debiased formulation of the identification result that, under a strong identification condition, enjoys a mixed-bias property, which means we can multiply estimation rates and make them insignificant to first order. Together with estimation results for a nuisance in this new debiased formulation, we obtain asymptotic normal estimates for long-term causal effects of the novel treatment, enabling inference.
Figure 1: Causal diagrams for surrogate settings with unobserved confounders. Dashed circles ($U$) indicate unobserved variables. Dotted circles ($Y$) indicate variables observed historically, but unobserved for novel treatments.
$A$$S$$U$$Y$
(a) A setting with unconfounded treatment ($A$) but confounded surrogate ($S$) and outcome ($Y$).
[…]
(b) The general setting we tackle with two kinds of possible short-term observations $S=(S_{1},S_{2})$. The identity of the components $S_{1}$ and $S_{2}$ within $S$ need not be known.
3.3 Causal Model
Up until now we worked only with an observed-data distribution and posed Eq. 1 as a statistical problem. We now introduce a causal model for the surrogate setting we study, for which we will show how the NPIV problem Eq. 1 helps identify a causal effect.
Potential outcomes.
We state our causal model in general terms using potential outcomes. The model is summarized in Fig. 1(b), which represents a causal diagram that is consistent with these assumptions. We assume the existence of variables $\widetilde{A},U_{1},S_{1}(a),U_{2}(a),S_{2}(u_{2}),Y(s_{1},s_{2},u_{2})$, and we abbreviate $S_{2}(a)=S_{2}(U_{2}(a))$, $Y(a)=Y(S_{1}(a),S_{2}(a),U_{2}(a))$.
Here we let $\widetilde{A}\in\{1,\dots,K+1\}$, where $\widetilde{A}\leq K$ denotes the historical experimental data and $\widetilde{A}=K+1$ denotes the novel treatment (we will return to data generation after the causal model). We are interested in inference on the average outcome under the novel treatment:
$\theta_{0}=\mathbb{E}[Y(K+1)].$
Our key assumptions characterizing the causal model are:
Assumption 1 (Unmeasured confounder and potential outcomes independence).
[…]
Assumption 2 (Randomization).
$A$ is independent of all potential outcomes and of $U_{1}$.
Assumption 3 (Existence of bridge function).
There exists $h^{\star}$ such that, for every $a\in[K]$, $\mathbb{E}[Y(a)-h^{\star}(S(a))]=0$.
Assumption 4 (Extrapolability).
For any $h:\mathcal{S}\to\mathbb{R}$
$(\mathbb{E}[Y(a)-h(S(a))]~{}\forall a\in[K])\implies\mathbb{E}[Y(K+1)-h(S(K+1))]=0.$ (6)
Assumption 1 captures the presence of confounding between short- and long-term outcomes. Assumption 2 captures that experiments are randomized so treatment assignment can be used as an IV.
Example.
It is straightforward to check that our causal assumptions hold under the following nonparametric structural causal model (SCM)
$Y=$ $f_{Y}(S_{1},S_{2},U_{2},\epsilon_{Y})+g_{Y}(U_{1},S_{2},U_{2},\epsilon_{Y})$
$S_{2}=$ $f_{S_{2}}(U_{2},\epsilon_{S_{2}})$
$S_{1}=$ $f_{S_{1}}(A,U_{1},\epsilon_{S_{1}})$
$U_{2}=$ $f_{U_{2}}(A,\epsilon_{U_{2}})$
$U_{1}=$ $f_{U_{1}}(\epsilon_{U_{1}})$
where $\epsilon_{Y}$, $\epsilon_{S_{2}}$, $\epsilon_{S_{1}}$, $\epsilon_{U_{2}}$, $\epsilon_{U_{1}}$ are independent. The key for assumption 3 to hold is that $U_{1}$ and $S_{1}$ enter the outcome equation in an additively separable manner. Aside from this additive restriction, this particular SCM is depicted in Fig. 1(b).
3.4 Data
Let $\widetilde{S}=(S_{1}(\widetilde{A}),S_{2}(\widetilde{A}))$ and $\widetilde{Y}=Y(\widetilde{A})$ be the factually observed outcomes.
Historical data set.
Let $A$ be uniformly distributed over $[K]$, let $Y,S$ be such that $(Y,S)\mid A=a\sim(\widetilde{Y},\widetilde{S})\mid\widetilde{A}=a$. Let $O=(A,S,Y)$. We have $N$ historical units for which we have observations $O_{i}=(A_{i},S_{i},Y_{i}),i=1,\ldots,N$ such that $O_{i}\sim O$. The treatment assignments $A_{1}\ldots,A_{N}$ of the historical units $i=1,\ldots,N$ is drawn following a completely randomized design such that each cell has exactly $n$ units. Any two $i\neq j$, $(S_{i},Y_{i})$, $(S_{j},Y_{j})$ are independent conditional on $A_{i},A_{j}$.
Novel data set.
Let $S^{(K+1)}\sim\widetilde{S}\mid\widetilde{A}=K+1$. We observe $n^{\prime}$ independent copies $S^{(K+1)}_{1},\ldots,S^{(K+1)}_{n^{\prime}}$ of $S^{(K+1)}$.
We now turn to the question of identification. Note that had we observed $Y^{(K+1)}\sim\widetilde{Y}\mid\widetilde{A}=K+1$ from the novel treatment, we would have trivially identified $\theta_{0}$ as $\mathbb{E}[Y^{(K+1)}]$ under Assumption 2. So, in the absence of outcome observations for the novel treatment, our goal is to combine the $N$ and $n^{\prime}$ datasets to identify $\theta_{0}$.
3.5 Surrogacy and set identification
It is immediate that under assumptions 3 and 4, $h^{\star}$ is a valid surrogate function in the sense that $E[Y(K+1)]=E[h^{\star}(S(K+1))]$.
That is, $h^{\star}$ connects causal effects on $Y$ to causal effects on $S$: if we could access $h^{\star}$, we would be able to know the average potential outcome of $Y$ under any treatment from the distribution of the average potential outcome of $S$ under that same intervention. We thus define the causal parameter of interest as $\theta^{\star}=\theta(h^{\star})$ where $\theta(h)=\mathbb{E}[h(S(a^{\prime}))]$ for the novel intervention $a^{\prime}$.
Set identification of $h^{\star}$ and point identification of $\theta_{0}$.
For any function $h:\mathcal{S}\to\mathbb{R}$, let $[\Pi h](a)=\mathbb{E}[h(S)\mid A=a]$. The following lemma characterizes $h^{\star}$ in terms of the data distribution in the historical cells.
Lemma 2 (Set identification of $h^{\star}$).
Under the causal assumptions 1-3, it holds that $h^{\star}\in\mathcal{H}_{0}$ where $\mathcal{H}_{0}$ is the set of functions $h:\mathcal{S}\to\mathbb{R}$ that satisfy the conditional moment restriction (CMR) $[\Pi h](a)=\mathbb{E}[Y\mid A=a]$ for every $a\in[K]$.
3.6 Point identification and strong identification
While $h^{\star}$ is not point identified a priori, it is not our parameter of interest: $\theta(h^{\star})$ is. We now present some assumptions under which $\theta(h^{\star})$ is point identified.
Definition 1 (Identification).
We say that $\theta(h^{\star})$ is point identified if for any solution $h$ of the CMR, that is for any $h\in\mathcal{H}_{0}$, $\theta(h)=\theta(h^{\star})$.
Let $\alpha$ be the importance sampling ratio between the distribution of $S$ in the target treatment arm ($A=a^{\prime}$) and in the historical experiments ($A\in[K]$), that is
$\alpha(s)=\frac{p_{S(K+1)}(s)}{\frac{1}{K}\sum_{a=1}^{K}p_{S(a)}(s)}=\frac{p_{S^{(K+1)}}(s)}{p_{S}(s)}.$ (7)
For any $h$ we can rewrite $\theta(h)$ as an expectation under the historical data experiment in terms of the importance sampling ratio:
$\theta(h)=\mathbb{E}[\alpha(S)h(S)].$ (8)
The function $\alpha$ is then the Riesz representer of the functional $\theta$. It is then immediate that $\theta(h^{\star})$ is point identified if, and only if, $\alpha\in\mathcal{N}(\Pi)^{\perp}=\overline{\mathcal{R}(\Pi^{*})}$.
Definition 2 (Strong identification).
We say that $\theta(h^{\star})$ is strongly identified if $\alpha\in\mathcal{R}(\Pi^{*}\Pi)$.
Note that strong identification trivially implies identification. Let $\Xi_{0}$ be the set of functions $\xi:\mathcal{S}\to\mathbb{R}$ such that $\alpha=\Pi^{*}\Pi\xi$. Strong identification is equivalent to $\Xi_{0}\neq\emptyset$. Under strong identification, $\theta(h^{\star})$ admits an additional representation, as we make explicit in the next lemma.
Lemma 3 (Double robustness, mixed bias).
[…]
$\psi(h,\xi)=h(S^{(K+1)})+\xi(S^{\prime})(Y-h(S)).$ (9)
Then it holds that
$\left\lvert\mathbb{E}[\psi(h,\xi)]-\theta^{\star}\right\rvert\leq\left\lVert\Pi(\xi-\xi_{0})\right\rVert\times\left\lVert\Pi(h-h_{0})\right\rVert,$ (10)
[…]
Remark 1.
The requirement that $S^{\prime}$ is an independent copy of $S$ will be easily satisfied at estimation time by splitting treatment arms in non-overlapping folds and drawing $S^{\prime}$ for the same cell as $(S,Y)$ but from a different fold.
Remark 2.
Unlike previous works, we get a mixed bias property in terms of weak norms of both $h$ and $\xi$. This is very specific to our setting though, as the key to this property is that we can easily draw an independent copy $S^{\prime}$ of $S$, which we can do here due to (i) the fact that we can group observations by value of $A$ (that is in treatment arms) and split these groups in folds, and (ii) the non-overlapping nature of the variables $A$, $S$, $Y$.
We now turn to the key result of this section.
Theorem 4 (Identification implies strong identification).
In the setting we consider in this section, identification and strong identification are equivalent, that is
$\overline{\mathcal{R}(\Pi^{*})}=\mathcal{R}(\Pi^{*})=\mathcal{R}(\Pi^{*}\Pi).$ (11)
4 Estimation of the debiasing nuisance function
Risk for debiasing nuisance.
Observe that $\xi\in\Xi_{0}$, that is $\Pi^{*}\Pi\xi=\alpha$, is the first order condition in the minimization w.r.t. functions $L_{2}(\mathcal{S})$ of the risk
$R_{2}(\xi)=$ $\frac{1}{2}\mathbb{E}\left[[\Pi\xi](A)^{2}\right]-\mathbb{E}\left[\alpha(S)\xi(S)\right]$ (12)
$=$ $\frac{1}{2}\mathbb{E}\left[[\Pi\xi](A)^{2}\right]-\mathbb{E}\left[\xi(S^{(K+1)})\right].$ (13)
Similarly to the case of the primary nuisance function $h$, we introduce a penalized cross-fold empirical analog of the population debiasing risk. Specifically, we let
$\widehat{R}_{2,\mu}(\xi)=\frac{1}{2K}\sum_{a=1}^{K}[\widehat{\Pi}_{0}\xi](a)[\widehat{\Pi}_{1}\xi](a)-\frac{1}{n^{\prime}}\sum_{A_{i}=a^{\prime}}\xi(S^{(K+1)}_{i})+\mu\|\xi\|_{2,N}^{2}$ (14)
Let $\widehat{\xi}$ be a minimizer of $\widehat{R}_{2,\mu}(\xi)$, and let $\xi_{0}$ be the minimum norm element of $\Xi_{0}\cap\mathcal{H}$. Our convergence rate results require the following condition.
Assumption 5 (Realizability for $\xi$).
It holds that $\Xi_{0}\cap\mathcal{H}\neq 0$.
Slow rate.
Let $\mathcal{R}^{\prime}_{n^{\prime}}(\mathcal{H},\delta)=E[\sup_{\xi\in\mathcal{H}:\|\xi\|_{(K+1)}\leq\delta}\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\epsilon_{i}\xi(S^{\prime}_{i})]$ be the localized Rademacher complexity of $\mathcal{H}$ on the new dataset ($A=K+1$). Let $(r^{\prime}_{n^{\prime}})^{2}$ be an upper bound on $\mathcal{R}^{\prime}_{n^{\prime}}(\mathcal{H},1)$.
Without any further assumptions, we obtain the following slow rate result.
Theorem 5 (Slow rate for $\xi$).
Let $\zeta>0$. Suppose that assumption 5 holds. Suppose that $\sqrt{\log(K/\zeta)}=o(\widetilde{\delta}_{n})$, and $\sqrt{\log\log\delta_{N}/N}=o(\delta_{N})$. Then it holds with probability $1-O(\zeta)$ that
$\|\Pi(\xi-\xi_{0})\|=\widetilde{O}(\delta_{N}^{1/2}+r^{\prime}_{n^{\prime}}).$ (15)
Fast rate.
Let $\delta^{\prime}_{n^{\prime}}$ be a solution to $\mathcal{R}^{\prime}_{n^{\prime}}(\mathcal{H},\delta^{\prime})\leq(\delta^{\prime}_{n^{\prime}})^{2}$. We can obtain a fast rate under the following so-called source condition on the Riesz representer $\alpha$.
Definition 3 (Source condition).
We say that the source condition with parameter $\beta$ holds for $\alpha$ if $\alpha=(\Pi^{*}\Pi)^{\beta}w_{0}$ for some $w_{0}:\mathcal{S}\to\mathbb{R}$.
Theorem 6 (Fast rate).
Let $\zeta>0$.Suppose that assumption 5 holds. Suppose that $\sqrt{\log(K/\zeta)}=o(\widetilde{\delta}_{n})$, and $\sqrt{\log\log\delta_{N}/N}=o(\delta_{N})$. Further suppose that the source condition holds with parameter $\beta\in(1,2]$ and that $w_{0}$ in the source condition is such that $\sup_{K}\|w_{0}\|<\infty$. Set $\lambda=\delta_{n^{\prime}}^{\frac{2}{2+\beta}}$. Then it holds with probability $1-O(\zeta)$ that
$\|\Pi(\widehat{\xi}-\xi_{0})\|=\widetilde{O}((\delta^{\prime}_{n^{\prime}})^{-\frac{1+\beta}{2+\beta}}).$ (16)
and
$\|\xi-\xi_{0}\|=\widetilde{O}((\delta^{\prime}_{n^{\prime}})^{-\frac{\beta}{2+\beta}}).$ (17)
Note that the convergence rates above are in terms of the norm induced by historical data distribution.
5 Semiparametric cross-fold estimator
Divide the historical data set in 4 folds $v=0,1,2,3$. Estimate $\widehat{h}$ on the first two folds and $\widehat{\xi}$ on the first two folds and the new treatment arm. For every index $i$ in fold $2$, let $j(i)$ be a corresponding observation index in fold 3, such that $i\mapsto j(i)$ is a one-to-one mapping, and let $(S_{i}^{\prime},Y_{i}^{\prime})=(S_{j(i)},Y_{j(i)})$111this requires that the two fold have the same number of observations.
$\widehat{\theta}=\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\widehat{h}(S_{i}^{(K+1)})+\frac{4}{N}\sum_{i}\xi(S_{i})(Y_{i}^{\prime}-\widehat{h}(S_{i}^{\prime})).$ (18)
Theorem 7.
Suppose that $\|\Pi(\widehat{h}-h_{0})\|\|\Pi(\widehat{\xi}-\xi_{0})\|=o_{P}((n^{\prime}\wedge N)^{-1/2})$ and that $\|\widehat{h}-h_{0}\|=o_{P}(1)$ and $\|\widehat{\xi}-\xi_{0}\|=o_{P}(1)$ Then
[…]
where $\sigma_{1}^{2}=\mathrm{Var}(h_{0}(S)\mid A=a^{\prime})$ and $\sigma_{2}^{2}=\mathrm{Var}(\xi_{0}(S)\mid A\in[K])\times\mathrm{Var}(Y-h_{0}(S))\mid A\in[K])$.
6 Numerical Experiments
We perform simulation experiments to study the behavior of cross-fold npJIVE. To begin, we fix a data-generating process with treatment effects $\mu_{a}$ generated from a normal distribution with variance $\sigma_{a}$, where the latter choice controls the strength of the instruments. The short-term outcome $S$ is discrete and generated from a multinomial distribution with probability vector drawn $\text{Dirichlet}(\mathbf{1}\cdot\exp(\mu_{a})$. The long-term outcome $Y$ is generated as a combination of non-linear function $h(S)=S+5\times(S>3)$, an unobserved confounder $U\sim\mathcal{N}(0,1)$, and standard normal noise. In the ‘historical experiments’ dataset $\mathfrak{H}$, we have $A=[1,\dots k]$ experiments that contain $n=50$ observations each that follows the above DGP. In a ‘target experiments’ dataset $\mathfrak{T}$, we have $n=100$ observations from the same DGP.
To begin, we compare the performance of a kernel-ridge implementation of $\hat{h}(S)$ trained on the historical data $\mathfrak{H}$ in recovering the true $h_{0}(S)$ in the target data $\mathfrak{T}$ using the weak norm $\frac{1}{K}\sum_{a}\hat{\mathbb{E}}(Y-\hat{h}(S)\mid A=a]^{2}/\sigma_{a}^{2}$ varying instrument strength $\sigma_{a}$ and number of instruments $k$. We report this weak norm in figure 2, and find that the cross-fold moments estimator empirically converges in the weak norm, with faster convergence when instruments are stronger.
Figure 2: Weak norm normalized by instrument strength, that is $\frac{1}{K}\sum_{a}\hat{\mathbb{E}}(Y-\hat{h}(S)\mid A=a]^{2}/\sigma_{a}^{2}$ varying instrument strength $\sigma_{a}$ and number of instruments evaluated on target experiments
Next, we perform a simulation experiments to compare the performance of NPJIVE with its non-cross-fold counterpart, the sieve minimum distance (SMD) estimator (Ai and Chen, 2003). Both estimators of $\hat{h}(S)$ involve fitting a kernel-ridge regression of the form $\hat{\beta}_{h}=(\Phi Q\Phi+\lambda I_{n})^{-1}\Phi QY$ on the historical experiments data, where $\Phi$ is the Gram matrix of the short-term outcomes $S$, and $Q$ is sparse selector matrix of dimensions $nK\times nK$ (where $n$ is the sample size in each experiment and $K$ is the number of instruments). The sparsity of $Q$ distinguishes our estimator from 2SLS/SMD: without cross-folds, $Q$ is block-diagonal with $n\times n$ matrices of ones on the diagonal, while the use of cross-folds introduces zeros to compute held-out averages (with maximal sparsity with two folds and minimal with $n-1$ folds, which is the conventional leave-one-out jackknife), as illustrated in fig 4.
Similar to above, we fix a DGP with treatment effects $\mu_{a}$ drawn from a normal distribution with variance $\sigma_{a}$, which governs instrument strength, and the short term outcome is generated from a uniform distribution $\mathsf{Unif}[\mu_{a}-0.5,\mu_{a}+0.5]$, and the long-term outcome is generated as a combination of a non-linear function $h(S)=S+0.5(S>0.25)-\sin(S)$ and an unobserved confounder $U$. In the ‘historical experiments’ dataset $\mathfrak{H}$, we have $A=[1,\dots,K=5000]$ experiments, each with $n$ observations each, where we shall vary $n$ to ‘very small’ and ‘moderate’ values. We compare the performance of SMD and npJIVE in recovering $\widehat{S}$ in the weak norm. We report results for two levels of instrument strength $\sigma_{a}$, and experiment size $n$, in figure 4, where the x-axis reports the bandwidth of the kernel (with the optimal value denoted by a dotted line). We find that a tuned npJIVE consistently weakly outperforms SMD, with large gaps with small weak experiments.
Figure 3: Q matrices for KRR implementation of SMD and npJIVE. Each matrix is $nK\times nK$, with black cells corresponding with zeroes and light cells corresponding with ones.
l Refer to caption
Figure 4: Weak norm of npJIVE and SMD, varying RBF kernel (inverse)bandwidth $\gamma$, instrument strength $\sigma_{a}$ (column) and sample size in each experiment (rows), with total number of instruments fixed at $5000$. We see that a tuned (i.e. optimal bandwidth) npJIVE consistently outperforms SMD, with gaps being largest when instruments are weak and experiments are small
7 Concluding remarks
The proliferation of experimentation on digital platforms provides a bountiful source of instruments (Peysakhovich and Eckles, 2018). This holds great potential to learn much more than what each such experiment taught us about the treatment that was experimented with. In particular, using it to understand the relationship between short- and long-term metrics (Bibaut et al., 2024), we may even be able to predict the effects of new experiments before observing any long-term outcome, even when the perfect surrogacy needed for observational surrogate index methods (Athey et al., 2019; Chen and Ritzwoller, 2023) fails. A key challenge is that, while we have many experiments, each experiment only tells us so little. Here we tackled the combined challenges of NPIV, many weak IVs, and inference on low dimensional functionals, our solution thereto newly enables statistical inference on long-term causal effects with nonparametric surrogate indices learned from many weak historical experiments.
[…]
Appendix A Proof of theorem 4
Proof.
That $\overline{\mathcal{R}(\Pi^{*})}=\mathcal{R}(\Pi^{*})$ directly follows from the fact that $\mathcal{R}(\Pi^{*})$ is a finite dimensional linear subspace, and therefore equals its closure.
Let $\phi_{a}:s\mapsto p[s\mid a]$. Let $r=\mathrm{rank}(\{\phi_{a}:a\in[K]\})$. There exists $a_{1},\ldots,a_{r}$ such that $\phi_{a_{1}},\ldots,\phi_{a_{r}}$ are linearly independent.
From now on, we proceed in four steps.
Step 1: $\Pi\phi_{a_{1}},\ldots,\Pi\phi_{a_{r}}$ are linearly independent.
Let $\gamma_{1},\ldots,\gamma_{r}$ be such that $\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}=0$. In particular must have that, for every $j=1,\ldots,r$, $\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}(a_{j})=0$, and therefore, by summation, $\sum_{1\leq i,j\leq r}\gamma_{i}\gamma_{j}\Pi\phi_{a_{i}}(a_{j})=0$, that is, by definition of $\Pi$,
$\sum_{1\leq i,j\leq r}\gamma_{i}\gamma_{j}\langle\phi_{a_{i}},\phi_{a_{j}}\rangle_{L_{2}(\mathcal{S})}=\langle\sum_{i=1}^{r}\gamma_{i}\phi_{a_{i}},\sum_{i=1}^{r}\gamma_{i}\phi_{a_{i}}\rangle_{L_{2}(\mathcal{S})}=0.$ (20)
Therefore, from positive definiteness, $\sum_{i=1}^{r}\gamma_{i}\phi_{a_{i}}$, which implies that $\gamma_{1}=\ldots=\gamma_{r}=0$, since $\phi_{a_{1}},\ldots,\phi_{a_{r}}$ are linearly independent. We have thus shown that $\Pi\phi_{a_{1}},\ldots,\Pi\phi_{a_{r}}$ are linearly independent.
Step 2: $\Pi^{*}\Pi\phi_{a_{1}},\ldots,\Pi^{*}\Pi\phi_{a_{r}}$ are linearly independent.
Let $\gamma_{1},\ldots,\gamma_{r}$ be such that $\sum_{i=1}^{r}\gamma_{i}\Pi^{*}\Pi\phi_{a_{i}}=0$. We must then have for every $j=1,\ldots,r,$ that
$\langle\sum_{i=1}^{r}\gamma_{i}\Pi^{*}\Pi\phi_{a_{i}},\phi_{a_{j}}\rangle_{L_{2}(\mathcal{S})}=0,$ (21)
and therefore, by summation, that
$\langle\sum_{i=1}^{r}\gamma_{i}\Pi^{*}\Pi\phi_{a_{i}},\sum_{i=1}^{r}\gamma_{i}\phi_{a_{j}}\rangle_{L_{2}(\mathcal{S})}=0,$ (22)
and thus, by definition of the adjoint, that
$\langle\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}},\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}\rangle=0.$ (23)
By positive definiteness, $\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}=0$, which implies that $\gamma_{1}=\ldots=\gamma_{r}=0$ since $\Pi\phi_{a_{1}},\ldots,\Pi\phi_{a_{r}}$ are linearly independent.
Step 3: $\mathrm{rank}(\Pi^{*})\leq r$.
We have that $\mathcal{R}(\Pi^{*})=\mathrm{Span}(s\mapsto p[a\mid s]:a\in[K])$. From Bayes formula,
$\mathcal{R}(\Pi^{*})=\mathrm{Span}\left(s\mapsto\frac{\phi_{a}(s)}{Kp(s)}:a\in[K]\right)\subseteq\mathrm{Span}\left(s\mapsto\frac{\phi_{a_{i}}(s)}{Kp(s)}:i=1,\ldots,r\right),$ (24)
since $\phi_{a_{1}},\ldots,\phi_{a_{r}}$ is a basis of $\mathrm{Span}(\phi_{a}:a\in[K])$.
Concluding.
We have that $\mathcal{R}(\Pi^{*}\Pi)\subseteq\mathcal{R}(\Pi^{*})$. From step 2, $\mathrm{rank}(\Pi^{*}Pi)\geq r$, and from step 3, $\mathrm{rank}(\Pi^{*})\leq r$. Therefore, we must have $\mathcal{R}(\Pi^{*}\Pi)=\mathcal{R}(\Pi^{*})$, which concludes the proof.
Appendix B Primary nuisance estimation proofs
For any two functions $f_{1},f_{2}:[K]\to\mathbb{R}$, let
$\langle f_{1},f_{2}\rangle=\frac{1}{K}\sum_{a=1}^{K}f_{1}(a)f_{2}(a).$ (25)
B.1 Local maximal inequality
Let $O_{1},\ldots,O_{N}\in\mathcal{O}$ be $N$ independent random variables. For any function $f:\mathcal{O}\to\mathbb{R}$, let
$\|f\|=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[f(O_{i})^{2}]}.$ (26)
Lemma 8 (Local maximal inequality).
Let $\mathcal{F}$ be a star-shaped class of functions such that $\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq 1$ and $\|f\|\leq r$. Suppose that $N^{-1/2}\sqrt{\log\log\delta_{N}}=o(\delta_{N})$. Let $b_{1},\ldots,b_{N}$ be fixed real numbers such that $\max_{i=1,\ldots,N}|b_{i}|\leq B$. Then, for any $u>0$, it holds with probability $1-e^{-u^{2}}$ that for every $f\in\mathcal{F}$
$\frac{1}{N}\sum_{i=1}^{N}b_{i}(f(O_{i})-\mathbb{E}[f(O_{i})])\lesssim B\left(\delta_{N}\|f\|+\delta_{N}^{2}+\frac{u\|f\|}{\sqrt{N}}+\frac{u^{2}}{N}\right).$ (27)
Proof.
Let $\mathcal{S}_{m}=\{f:\|f\|\leq r_{m}\}$ where $r_{m}=2^{m}\delta_{N}$. From Talagrand’s or alternatively Bousquet’s inequality, it holds with probability $1-e^{-u^{2}}$ that for every $f\in\mathcal{S}_{m}$,
$\frac{1}{N}\sum_{i=1}^{N}b_{i}(f(O_{i})-\mathbb{E}[f(O_{i})])\lesssim\mathbb{E}\left[\sup_{f\in\mathcal{S}_{m}}\frac{1}{N}\sum_{i=1}^{N}\epsilon_{i}b_{i}f(O_{i})\right]+B(\frac{ur_{m}}{\sqrt{N}}+\frac{u^{2}}{N}).$ (28)
From the contraction lemma for Rademacher processes,
$\mathbb{E}\left[\sup_{f\in\mathcal{S}_{m}}\frac{1}{N}\sum_{i=1}^{N}\epsilon_{i}b_{i}f(O_{i})\right]\lesssim B\mathcal{R}_{N}(\mathcal{F},r_{m}).$ (29)
Let $m(f)=\min\{m\geq 0:f\in\mathcal{S}_{m}\}$. From a union bound, with probability $1-e^{-u^{2}}$, for every $f\in\mathcal{F}$,
$\frac{1}{N}\sum_{i=1}^{N}b_{i}(f(O_{i})-\mathbb{E}[f(O_{i})])\lesssim B\left(\mathcal{R}_{N}(\mathcal{F},r_{m(f)})+(u+\sqrt{\log\log_{2}\delta_{N}})\frac{r_{m(f)}}{\sqrt{N}}+\frac{u^{2}\log\log_{2}\delta_{N}}{N}\right).$ (30)
If $m(f)=0$, $\mathcal{R}_{N}(\mathcal{F},r_{m(f)})=\mathcal{R}_{N}(\mathcal{F},\delta_{N})\leq\delta_{N}^{2}$ by definition of the critical radius. If $m(f)\geq 1$,
$\frac{\mathcal{R}_{N}(\mathcal{F},r_{m(f)})}{r_{m(f)}}\leq\delta_{N},$ (31)
therefore $\mathcal{R}_{N}(\mathcal{F},r_{m(f)})\leq\delta_{N}\|f\|$. Therefore, with probability $1-e^{-u^{2}}$, it holds that for every $f\in\mathcal{F}$,
$\frac{1}{N}\sum_{i=1}^{N}b_{i}(f(O_{i})-\mathbb{E}[f(O_{i})])\lesssim B\left(\delta_{N}\|f\|+\delta_{N}^{2}+\frac{u\|f\|}{\sqrt{N}}+\frac{u^{2}}{N}\right).$ (32)
B.2 Decomposition of the difference between population and empirical risk
We have that
$2(\widehat{R}_{1,0}(h)-R_{1}(h))$ (33)
$=$ $\langle\widehat{\Pi}_{0}(h_{0}-h+\eta),\widehat{\Pi}_{1}(h_{0}-h+\eta)\rangle-\langle\Pi(h_{0}-h),\Pi(h_{0}-h)\rangle$ (34)
$=$ $\langle(\widehat{\Pi}_{0}-\Pi)(h-h_{0}),(\widehat{\Pi}_{1}-\Pi)(h-h_{0})\rangle$ (35)
$+\sum_{v=0,1}\langle\Pi(h-h_{0}),(\widehat{\Pi}_{v}-\Pi)(h-h_{0})\rangle$ (36)
$+\sum_{v=0,1}\langle\Pi(h-h_{0}),(\widehat{\Pi}_{v}-\Pi)\eta\rangle$ (37)
$+\sum_{v=0,1}\langle(\widehat{\Pi}_{1-v}-\Pi)(h-h_{0}),(\widehat{\Pi}_{v}-\Pi)\eta\rangle$ (38)
$+\langle(\widehat{\Pi}_{0}-\Pi)\eta,(\widehat{\Pi}_{1}-\Pi)\eta\rangle$ (39)
B.3 U-process terms
Denote $\mathcal{G}=\mathcal{H}-h_{0}$. For any $g\in\mathcal{G}$, let $g^{0}=g-\Pi g$, and let $\mathcal{G}^{0}=\{g-\Pi g:g\in\mathcal{G}\}$ For any such $g^{0}$, let
$Z_{N,1}^{0}(g)=\langle\widehat{\Pi}_{0}g^{0},\widehat{\Pi}_{1}g^{0}\rangle.$ (40)
For any $a\in[K]$, $\delta>0$ let
$\mathcal{R}_{n,a}(\mathcal{G}^{0},\delta)=\mathbb{E}\left[\sup_{g^{0}\in\mathcal{G}^{0}:\|g^{0}\|_{2,a}\leq\delta}\frac{2}{n}\sum_{i:A_{i}=a,V_{i}=0}\epsilon_{i}g^{0}(S_{i})\right].$ (41)
Lemma 9.
[…]
$Z_{N}^{0}(g)=O\left(\left(\widehat{\delta}_{N}\widetilde{\delta}_{n}\|g^{0}\|^{2}+\widehat{\delta}_{N}\widetilde{\delta}_{n}^{2}\|g^{0}\|+\widehat{\delta}_{N}^{2})\right)\wedge\widehat{r}_{N}^{2}\right),$ (42)
where the constant in $O(\cdot)$ is uniform over $\mathcal{G}^{0}$.
Proof.
From lemma 8, a union bound and the condition $\sqrt{\log(\zeta/K)/n}=o(\widetilde{\delta}_{n})$, it holds with probability $1-\zeta$ that, for every $a\in[K]$,
$[\widehat{\Pi}_{0}g^{0}](a)=O(\widetilde{\delta}_{n}\|g^{0}\|_{2,a}+\widetilde{\delta}_{n}^{2}).$ (43)
Let $\mathcal{E}$ be the event that the above display holds. Conditional on $\mathcal{D}_{0}$, under thes assumption that $\sqrt{\log\log\widetilde{\delta}_{n}/n}=o(\widetilde{\delta}_{n})$ and $\widehat{\delta}_{N}=o(\sqrt{\log\log\widehat{\delta}_{N}/N})$, from lemma 8, it holds that with probability $1-\zeta$ that, for any $g^{0}\in\mathcal{G}^{0}$,
$Z_{N,1}(g^{0})\leq\widehat{\delta}_{N}\|\widehat{\Pi}_{0}g^{0}\times g\|_{|\mathcal{D}_{0}}+\widehat{\delta}_{N}^{2}.$ (44)
[…]
Under the event $\mathcal{E}$, for any $g^{0}\in\mathcal{G}^{0}$,
$\|\widehat{\Pi}_{0}g^{0}\times g\|_{|\mathcal{D}_{0}}=$ $\sqrt{\frac{2}{N}\sum_{i:V_{i}=1}\left\lvert\widehat{\Pi}_{0}g^{0}(A_{i})\right\rvert^{2}\mathbb{E}[(g^{0}(S_{i}))^{2}]}$ (45)
$\leq$ $\sqrt{\frac{2}{N}\sum_{i:V_{i}=1}(\widetilde{\delta}_{n}\|g^{0}\|_{2,A_{i}}+\widetilde{\delta}_{n}^{2})^{2}\|g^{0}\|_{2,A_{i}}^{2}}$ (46)
$\lesssim$ $\widetilde{\delta}_{n}\sqrt{\frac{2}{N}\sum_{i:V_{i}=1}\|g^{0}\|_{2,A_{i}}^{4}}+\widetilde{\delta}_{n}^{2}\|g^{0}\|$ (47)
$\lesssim$ $\widetilde{\delta}_{n}\|g^{0}\|^{2}+\widetilde{\delta}_{n}^{2}\|g^{0}\|.$ (48)
A $(1-\zeta)$-probability $O(\widehat{r}_{N}^{2})$ bound directly follows from applying Talagrand’s or Bousquet’s inequality conditionally on $\mathcal{D}_{0}$. The claim then follows from a union bound. ∎
Now, let
$Z_{N,2}^{0}(g)=\langle\widehat{\Pi}_{0}\eta,\widehat{\Pi}_{1}g^{0}\rangle$ (49)
Lemma 10.
Let $\widetilde{\delta}_{N}$ be an upper bound on the critical radius of $\mathcal{G}^{0}$. Let $\zeta\in(0,1)$. Suppose that $\log(1/\zeta)=o(\log K)$, $\sqrt{\log(1/\zeta)/N}=o(\widetilde{\delta}_{N})$ and $\sqrt{\log\log\widetilde{\delta}_{N}/N}=o(\widetilde{\delta}_{N})$. Then, with probability $1-O(\zeta)$, for every $g\in\mathcal{G}$,
$Z_{N,2}^{0}(g)=O\left(\sqrt{\frac{\log K}{n}}(\widetilde{\delta}_{N}\|g^{0}\|+\widetilde{\delta}_{N}^{2})\right).$ (50)
Proof.
From Hoeffding, a union bound, and that $\log(1/\zeta)=o(\log K)$, it holds with probability $1-\zeta$ that
$\max_{a\in[K]}|[\widehat{\Pi}_{0}\eta](a)|=O\left(\sqrt{\frac{\log K}{n}}\right).$ (51)
Observe that
$Z_{N,2}^{0}(g)=\frac{1}{N/2}\sum_{i:V_{i}=1}[\widehat{\Pi}_{0}\eta](A_{i})g^{0}(S_{i}).$ (52)
The conclusion then follows by applying lemma 8 conditional on $\mathcal{D}_{0}$, on the event that the above display holds, with $b_{i}=[\widehat{\Pi}_{0}\eta](A_{i})|$. ∎
B.4 Empirical process terms
B.4.1 First empirical process term
We have that
$\langle\Pi(h-h_{0}),(\widehat{\Pi}_{1}-\Pi)(h-h_{0})\rangle$ (53)
$=\frac{2}{N}\sum_{i:V_{i}=1}[\Pi(h-h_{0})](A_{i})(h-h_{0})(S_{i})-E[[\Pi(h-h_{0})](A_{i})(h-h_{0})(S_{i})],$ (54)
Therefore, from a direct application of lemma 8, we have that, as long as $\sqrt{\log(1/\zeta)/N}=o(\bar{\delta}_{N})$ and $\sqrt{\log\log(\delta_{N})/N}=o(\delta_{N})$, it holds with probability $1-2\zeta$ that for every $h\in\mathcal{H}$,
$\langle\Pi(h-h_{0}),(\widehat{\Pi}_{1}-\Pi)(h-h_{0})\rangle$ (55)
$\leq\left\lVert[\Pi(h-h_{0})]\times(h-h_{0})\right\rVert\bar{\delta}_{N}+\bar{\delta}_{N}^{2}$ (56)
$\leq\left\lVert[\Pi(h-h_{0})]\right\rVert\bar{\delta}_{N}+\bar{\delta}_{N}^{2}.$ (57)
B.4.2 Second empirical process term
Let $\breve{\delta}_{N}$ be an upper bound on the critical radius of $\Pi\mathcal{G}$.
We have that
$\langle\Pi g,\widehat{\Pi}_{0}\eta\rangle=$ $\frac{1}{N/2}\sum_{i:V_{i}=0}[\Pi g](A_{i})\eta_{i}-\mathbb{E}[[\Pi g](A_{i})\eta_{i}],$ (58)
Since the $\eta_{i}$’s are 1-subGaussian, it holds with probability $1-\zeta$ that
$\max_{i\in[N]}|\eta_{i}|=O(\sqrt{\log(N/\zeta)}).$ (59)
Therefore, from lemma 8 applied conditional on $\mathcal{D}_{0}$, on the event that the above display holds, it holds with probability $1-O(\zeta)$ that, provided the lemma’s conditions on $\zeta$ and $\log\log\breve{\delta}_{N}$ hold,
$\langle\Pi g,\widehat{\Pi}_{0}\eta\rangle=O\left(\sqrt{\log N}\left(\breve{\delta}_{N}\|\Pi g\|+\breve{\delta}_{N}^{2}\right)\right).$ (60)
B.5 Empirical and population risk difference for the primary nuisance
Term CR notation CR classes Rate (up to $\log$ terms)
$\langle(\widehat{\Pi}_{0}-\Pi)g,(\widehat{\Pi}_{1}-\Pi)g\rangle$ $\widetilde{\delta}_{n}$, $\widehat{\delta}_{N}$ $\mathcal{G}^{0}$, $(\widehat{\Pi}_{0}\mathcal{G}^{0})\times\mathcal{G}^{0}$ $\widehat{\delta}_{N}\widetilde{\delta}_{n}\|g^{0}\|^{2}+\widehat{\delta}_{N}\widetilde{\delta}_{n}^{2}\|g^{0}\|+\widehat{\delta}_{N}^{2}$
$\langle\Pi g,(\widehat{\Pi}_{v}-\Pi)g\rangle$ $\bar{\delta}_{N}$ $(\Pi\mathcal{G})\times\mathcal{G}$ $\bar{\delta}_{N}\|\Pi g\|+\bar{\delta}_{N}^{2}$
$\langle\Pi g,\widehat{\Pi}_{v}\eta\rangle$ $\breve{\delta}_{N}$ $\Pi\mathcal{G}$ $\breve{\delta}_{N}\|\Pi g\|+\breve{\delta}_{N}^{2}$
$\langle(\widehat{\Pi}_{1-v}-\Pi)g,(\widehat{\Pi}_{v}-\Pi)\eta\rangle$ $\widetilde{\delta}_{N}$ $\mathcal{G}^{0}$ $n^{-1/2}(\widetilde{\delta}_{N}\|g^{0}\|+\widetilde{\delta}_{N}^{2})$
Table 1: Summary of localized rates for the terms in the equicontinuity difference
Lemma 11.
Suppose the assumptions of the above lemmas hold. Suppose that $\widetilde{\delta}_{n}^{2}=O(n^{-1/2})$ and that $n^{-1/2}=O(\widetilde{\delta}_{n})$. Let $\delta_{N}=\max(\widetilde{\delta}_{N},\breve{\delta}_{N},\widehat{\delta}_{N})$. Then it holds with probability at least $1-O(\zeta)$ that for every $h$ in $\mathcal{H}$,
$\widehat{R}_{1,0}(h)-R_{1}(h)=O\left(\delta_{N}\|\Pi(h-h_{0})\|+\delta_{N}\widetilde{\delta}_{n}\|h-h_{0}\|^{2}+\delta_{N}n^{-1/2}\|h-h_{0}\|+\delta_{N}^{2}\right).$ (61)
B.6 Weak norm convergence rate for the primary nuisance
Lemma 12 (Convergence primary nuisance).
Suppose the assumptions of 11 hold, and that $\delta_{N}\widetilde{\delta}_{n}=O(\lambda)$. Then it holds with probability $1-O(\zeta)$ that
$\|\Pi(\widehat{h}-h_{0})\|^{2}=O(\lambda+\delta_{N}\delta_{n}).$ (62)
Proof.
All the statements in this proof hold with probability $1-ce^{-\delta}$. From lemma 11, optimality of $\widehat{h}$ for the empirical risk, and a another application of lemma 11,
$\|\Pi(\widehat{h}-h_{0})\|^{2}$ (63)
$\leq$ $\|\Pi(h^{*}-h_{0})\|^{2}$ (64)
$+O\left(\delta_{N}(\|\Pi(\widehat{h}-h^{*})\|+\|\Pi(h^{*}-h_{0})\|)\right.$ (65)
$\qquad+\delta_{N}\delta_{n}(\|\widehat{h}-h^{*}\|^{2}+\|h^{*}-h_{0}\|^{2})$ (66)
$\qquad+\delta_{N}n^{-1/2}(\|\widehat{h}-h^{*}\|+\|h^{*}-h_{0}\|)$ (67)
$\qquad+\left.\delta_{N}^{2}\right)$ (68)
$+\lambda(\|h^{*}\|_{2,N}^{2}-\|\widehat{h}\|_{2,N}^{2})$ (69)
$\leq$ $2\|\Pi(h_{0}-h^{*})\|^{2}$ (70)
$+O(\delta_{N}\|\Pi(\widehat{h}-h^{*})\|+\delta_{N}\delta_{n}\|\widehat{h}-h^{*}\|^{2}+\delta_{N}n^{-1/2}\|\widehat{h}-h^{*}\|+\delta_{N}\delta_{n})$ (71)
$+\lambda(\|h^{*}\|_{2,N}^{2}-\|\widehat{h}\|_{2,N}^{2}).$ (72)
From an analysis of the second-order polynomial $t\mapsto\|\Pi(h^{*}+t(\widehat{h}-h^{*})-h_{0})\|^{2}+\|h^{*}+t(\widehat{h}-h^{*})-h_{0}\|^{2}$,
$\|\Pi(\widehat{h}-h^{*})\|^{2}+\lambda\|\widehat{h}-h^{*}\|^{2}\leq\|\Pi(\widehat{h}-h_{0})\|^{2}-\|\Pi(h^{*}-h_{0})\|^{2}-\lambda(\|h^{*}\|^{2}-\|\widehat{h}\|^{2}).$ (73)
Combining this with the preceding display and the fact that $\|h^{*}\|^{2}-\|\widehat{h}\|^{2}-\|h^{*}\|^{2}_{2,N}-\|\widehat{h}\|^{2}_{2,N}=O(\|\widehat{h}-h^{*}\|\delta_{N}+\delta_{N}^{2})$ from lemma 8 yields that
$\|\Pi(\widehat{h}-h^{*})\|^{2}+\lambda\|\widehat{h}-h^{*}\|^{2}$ (74)
$\leq$ $\|\Pi(h^{*}-h_{0})\|^{2}$ (75)
$+O(\delta_{N}\|\Pi(h^{*}-h_{0})\|+\delta_{N}\delta_{n}\|\widehat{h}-h^{*}\|^{2}+\delta_{N}\delta_{n}^{2}\|\widehat{h}-h^{*}\|+\delta_{N}\delta_{n}+\lambda\delta_{N}\|\widehat{h}-h^{*}\|).$ (76)
Applying the AM-GM inequality to $\delta_{N}\|\Pi(h^{*}-h_{0})\|$ and using that $\delta_{N}\delta_{n}=O(\lambda)$ and that $\|\Pi(h^{*}-h_{0})\|^{2}=O(\lambda)$ yields that
$\frac{1}{2}\|\Pi(\widehat{h}-h^{*})\|^{2}+\frac{\lambda}{2}\|\widehat{h}-h^{*}\|^{2}=O(\lambda+\delta_{N}\delta_{n}),$ (77)
which yields the claim. ∎
B.7 Consistency in strong norm of the primary nuisance estimator
We have that
$\frac{1}{2}\|\Pi(\widehat{h}-h^{\dagger})\|^{2}\leq$ $R_{1}(\widehat{h})-R_{1}(h^{\dagger})$ (78)
$\leq$ $\widehat{R}_{1,0}(\widehat{h})-\widehat{R}_{1}(h^{\dagger})+O(\delta_{N}\|\Pi(\widehat{h}-h^{\dagger})\|+\delta_{N}\widetilde{\delta}_{n})$ (79)
$\leq$ $O(\delta_{N}\|\Pi(\widehat{h}-h^{\dagger})\|+\delta_{N}\widetilde{\delta}_{n})-\lambda(\|\widehat{h}\|_{2,N}^{2}-\|h^{\dagger}\|^{2}_{2,N}).$ (80)
Therefore, from AM-GM,
$\|\widehat{h}\|_{2,N}^{2}-\|h^{\dagger}\|^{2}_{2,N}=O\left(\frac{\delta_{N}\widetilde{\delta}_{n}}{\lambda}\right).$ (81)
Therefore, from the definition of $h^{\dagger}$, which implies the limit of the LHS is lower bounded by zero, we have that $\lim\|\widehat{h}\|=\|h^{\dagger}\|$, which implies the conclusion.
Appendix C Convergence of the debiasing nuisance
C.1 Regularization bias under source
Lemma 13.
Under the source condition for $\alpha$ with parameter $\beta$, we have that
$\|\xi-\xi_{0}\|=O(\|w_{0}\|^{2}\lambda^{\beta})\qquad\text{and}\qquad\|\Pi(\xi-\xi_{0})\|^{2}=O(\|w_{0}\|^{2}\lambda^{1+\beta}).$ (82)
The proof is identical to that of lemma 5 in Bennett et al. [2023b].
C.2 Decomposition of the equicontinuity term
The equicontinuity term decomposes as the sum of quadratic term and of a linear term as follows:
$(R(\xi)-R(\xi_{0}))-(\widehat{R}(\xi)-\widehat{R}(\xi_{0}))$ (83)
$=$ $Q(\xi,\xi_{0})+L(\xi,\xi_{0}),$ (84)
where
$Q(\xi,\xi_{0})$ (85)
$=$ $\frac{1}{2}\langle(\widehat{\Pi}_{0}-\Pi)(\xi-\xi_{0}),(\widehat{\Pi}_{1}-\Pi)(\xi-\xi_{0})\rangle$ (86)
$+\sum_{v=0,1}\left\{\langle\Pi(\xi-\xi_{0}),(\widehat{\Pi}_{v}-\Pi)(\xi-\xi_{0})\rangle\right.$ (87)
$\qquad+\langle(\widehat{\Pi}_{v}-\Pi)\xi_{0},(\widehat{\Pi}_{1-v}-\Pi)(\xi-\xi_{0})\rangle$ (88)
$\left.\qquad+\langle\Pi\xi_{0},(\widehat{\Pi}_{v}-\Pi)(\xi-\xi_{0})\rangle\right\}.$ (89)
and
$L(\xi,\xi_{0})=E[\xi(S^{\prime})-\xi_{0}(S^{\prime})].$ (90)
C.3 Bounding the equicontinuity term
Following the same arguments as for the primary nuisance, we obtain that
$Q(\xi,\xi_{0})=\widetilde{O}$ $\left(\widehat{\delta}_{N}\delta_{n}\|\xi-\xi_{0}\|^{2}+\widehat{\delta}_{N}\widetilde{\delta}_{n}^{2}\|\xi-\xi_{0}\|+\widehat{\delta}_{N}^{2}\right.$ (91)
$+\bar{\delta}_{N}\|\Pi(\xi-\xi_{0})\|+\bar{\delta}_{N}^{2}$ (92)
$+n^{-1/2}\widetilde{\delta}_{N}\|\xi-\xi_{0}\|+\widetilde{\delta}_{N}^{2}$ (93)
$+\left.\widetilde{\delta}_{N}\|\xi-\xi_{0}\|+\widetilde{\delta}_{N}^{2}\right)$ (94)
$=$ $\widetilde{O}\left(\delta_{N}\|\xi-\xi_{0}\|+\delta_{N}\widetilde{\delta}_{n}\|\xi-\xi_{0}\|^{2}+\delta_{N}^{2}\right).$ (95)
From lemma 8,
$L(\xi,\xi_{0})=\widetilde{O}\left(\delta^{\prime}_{n^{\prime}}\|\xi-\xi_{0}\|+(\delta_{n^{\prime}}^{\prime})^{2}\right).$ (96)
Therefore, since $n^{\prime}=o(N)$, we have that
$(R(\xi)-R(\xi_{0}))-(\widehat{R}(\xi)-\widehat{R}(\xi_{0}))=\widetilde{O}\left(\delta^{\prime}_{n^{\prime}}\|\xi-\xi_{0}\|+(\delta_{n^{\prime}}^{\prime})^{2}\right).$ (97)
C.4 Weak norm bound
From an analysis of $t\mapsto\|\Pi(t(\widehat{\xi}-\xi^{*})+(\xi^{*}-\xi_{0}))\|^{2}+\lambda\|(t(\widehat{\xi}-\xi^{*})+\xi^{*})\|^{2}$, it holds that
$\|\Pi(\widehat{\xi}-\xi^{*})\|^{2}+\lambda\|\widehat{\xi}-\xi^{*}\|\leq$ $\|\Pi(\widehat{\xi}-\xi_{0})\|^{2}-\|\Pi(\xi^{*}-\xi_{0})\|^{2}+\lambda\{\|\widehat{\xi}\|^{2}-\|\xi^{*}\|^{2}\}.$ (98)
From the equicontinuity term bound applied twice and optimality of $\widehat{\xi}$ for the penalized empirical risk, we have
$\|\Pi(\widehat{\xi}-\xi^{*})\|^{2}+\lambda\|\widehat{\xi}-\xi^{*}\|^{2}\leq$ $\|\Pi(\xi^{*}-\xi_{0})\|^{2}+\widetilde{O}\left(\delta_{n^{\prime}}^{\prime}\|\widehat{\xi}-\xi_{0}\|+(\delta_{n^{\prime}}^{{}^{\prime}})^{2}\right)$ (99)
$+\lambda\{(\|\widehat{\xi}\|^{2}-\|\widehat{\xi}\|_{2,N}^{2})-(\|\xi^{*}\|^{2}-\|\xi^{*}\|_{2,N}^{2})\}$ (100)
$\leq$ $\lambda^{1+\beta}+\widetilde{O}\left(\delta^{\prime}_{n^{\prime}}\|\widehat{\xi}-\xi^{*}\|+\delta_{n^{\prime}}^{\prime}\|\xi^{*}-\xi_{0}\|+(\delta_{n^{\prime}}^{\prime})^{2}\right)$ (101)
$\leq$ $\lambda^{1+\beta}+\widetilde{O}\left(\delta^{\prime}_{n^{\prime}}\|\widehat{\xi}-\xi^{*}\|+\delta_{n^{\prime}}^{\prime}\lambda^{\beta/2}+(\delta_{n^{\prime}}^{\prime})^{2}\right),$ (102)
where we have used in the second inequality that $\|\Pi(\xi^{*}-\xi_{0})\|^{2}\leq\lambda^{1+\beta}$ from lemma 13, and that $(\|\widehat{\xi}\|^{2}-\|\widehat{\xi}\|_{2,N}^{2})-(\|\xi^{*}\|^{2}-\|\xi^{*}\|_{2,N}^{2})=O(\delta_{N}\|\widehat{\xi}-\xi^{*}\|+\delta_{N}^{2})$ (see proof of theorem 4 in Bennett et al. [2023b]), and in the third inequality that $\|\xi^{*}-\xi_{0}\|^{2}\leq\lambda^{\beta}$, from lemma 13.
Therefore,
$\|\Pi(\widehat{\xi}-\xi^{*})\|^{2}=$ $\widetilde{O}\left(\sup_{x\in\mathbb{R}}-\lambda x^{2}+\delta_{n^{\prime}}^{\prime}x+\lambda^{1+\beta}+\delta_{n^{\prime}}^{\prime}\lambda^{\beta/2}+(\delta_{n^{\prime}}^{\prime})^{2}\right)$ (103)
$=$ $\widetilde{O}\left(\lambda^{-1}(\delta_{n^{\prime}}^{\prime})^{2}+\lambda^{1+\beta}+\delta_{n^{\prime}}^{\prime}\lambda^{\beta/2}\right),$ (104)
and then, for $\lambda=(\delta_{n^{\prime}}^{\prime})^{\frac{2}{2+\beta}}$, we have
$\|\Pi(\widehat{\xi}-\xi^{*})\|=\widetilde{O}\left((\delta_{n^{\prime}}^{\prime})^{\frac{1+\beta}{2+\beta}}\right),$ (105)
and since from lemma 13, we have $\|Pi(\xi^{*}-\xi_{0})\|=O(\lambda^{\frac{1+\beta}{2}})=O((\delta_{n^{\prime}}^{\prime})^{\frac{1+\beta}{2+\beta}})$, the triangle inequality implies the claim.
C.5 Strong norm bound
From (102),
$\|\widehat{\xi}-\xi^{*}\|^{2}=\widetilde{O}\left(\lambda^{\beta}+\frac{\delta_{n^{\prime}}^{\prime}}{\lambda}\|\widehat{\xi}-\xi^{*}\|+\delta_{n^{\prime}}^{\prime}\lambda^{\frac{\beta}{2}-1}\right).$ (106)
From the AM-GM inequality,
$\|\widehat{\xi}-\xi^{*}\|^{2}=\widetilde{O}\left(\lambda^{\beta}+\lambda^{-2}(\delta_{n^{\prime}}^{\prime})^{2}+\delta_{n^{\prime}}^{\prime}\lambda^{\frac{\beta}{2}-1}\right).$ (107)
For $\lambda=(\delta_{n^{\prime}}^{\prime})^{\frac{2}{2+\beta}}$, we then have that
$\|\widehat{\xi}-\xi^{*}\|^{2}=\widetilde{O}\left((\delta_{n^{\prime}}^{\prime})^{\frac{2\beta}{2+\beta}}\right),$ (108)
and since $\|\xi^{*}-\xi_{0}\|^{2}=O(\lambda^{\beta})=O((\delta_{n^{\prime}}^{\prime})^{\frac{2\beta}{2+\beta}})$, the triangle inequality implies the claim.